OpenStudy (anonymous):

(x^2-5x)/5x=x-1 is this correct?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

no

OpenStudy (anonymous):

also no

OpenStudy (anonymous):

\[\frac{x^2-5x}{5x}=\frac{x(x-5)}{5x}=\frac{x-5}{5}\]

OpenStudy (anonymous):

can cancel the x but not the 5 since 5 is not a factor of the first term.

OpenStudy (anonymous):

oh yeah-thanks

OpenStudy (anonymous):

In general it's not true but it is for x=0.

OpenStudy (anonymous):

no its not

OpenStudy (anonymous):

not in the form its in ,

OpenStudy (anonymous):

Form is temporary. Changing the form doesn't change the solution set, as long as you're careful with zero divisors. Here it's exactly the same as x/5 - 1 = x - 1. Which has solution x=0.

OpenStudy (anonymous):

dont worry :|

OpenStudy (anonymous):

what on earth?

OpenStudy (anonymous):

remember, they are different expressions

OpenStudy (anonymous):

its like saying graph \[y=2x-2\] and graph \[\frac{y}{2x-2} =1\]

OpenStudy (anonymous):

one has a hole at x=1

OpenStudy (anonymous):

Interesting point. I guess the subtlety lies in the assumption that we never have divisors of zero, and I broke that assumption. I really should know better lol.